# A new twist on the Majorana surface code: Bosonic and fermionic defects for fault-tolerant quantum computation

### v3

@@ IV.1 Code Deformation
IV.1.1 Moving Bosonic Twists
@@
IV.1.2 Moving Fermionic Twists
@@
IV.2 State Preparation of BT Qubits and Clifford Gates via Braiding Twists
@@
We will now discuss how to prepare logical Pauli eigenstates and perform fault-tolerant Clifford gates using BT qubits. The latter will be achieved purely by braiding $Z_{2}$ and $Z_{3}$ twists, in analogy to Clifford gate procedures in the bosonic Color Code [Bombin_twists_code_deformation] (though Clifford gates may also be achieved through lattice surgery; see Section IV.5). Our MSC scheme requires no ancilla logical qubits for the implementation of single-qubit Clifford gates, unlike other approaches [MjFerm_Surface_Code, Roadmap_to_MSCs]. We shall consider BT qubits stored sparsely in quartets of either $Z_{2}$ (storing one qubit in four twists) or $Z_{3}$ (two qubits in four twists) bosonic twists. We will see that we need only consider two sub-types of $Z_{2}$ twist, which we take (without loss of generality) to be $T_{\mathsf{B}}$ and $T_{\mathsf{R}}$.
To prepare four $Z_{2}$ twists in an eigenstate of a given logical Pauli operator, we start with a MSC patch hosting these twists, such as the one of Figure 1. For example, to prepare this patch in the $+1$ eigenstate of $\bar{X}$ (a red string operator), we initialize the patch in the $+1$ eigenstate of all red bond operators. This prepares $\bar{X}$ and green and blue plaquette operators in their $+1$ eigenstates since they are all products of red bond operators. To prepare red plaquette operators in their $+1$ eigenstates, we measure them and apply red string operators that take the state to the code space given the observed syndrome (in practice, we would not apply this operator and instead, redefine the stabilizers). Since $\bar{X}$ and the green/blue plaquette operators commute with the red plaquette operators, these measurements do not change their eigenvalues from the preceding step. The inequivalent choices of string in the last step differ by $\bar{X}$; this is again immaterial given we already had a $+1$ eigenstate of $\bar{X}$. This procedure thus prepares the patch in the $+1$ eigenstate of $\bar{X}$ in the code space. We can then move the twists away from the boundaries of the patch using code deformation (cf. Section IV.1). Finally, we can fuse the patch with the rest of the lattice hosting logical information by adding stabilizer operators that bridge the gap between the boundaries of the two lattices.
@@
IV.3 State Preparation of FT Qubits and Logical Majorana Parity Measurements via Lattice Surgery
@@
IV.3.1 Bilinear Measurement and Eigenstate Preparation
@@
Stabilizing Logical Operators
@@
IV.3.2 Arbitrary Fermion Parity Measurement
@@
To explain the effect of these $i\Upsilon_{j}\tilde{\Upsilon}_{j}$ measurements, we consider the example of Figure 12 more closely. Here we will use $\Gamma$ to denote the total fermion parity operator of the main lattice (containing $\Upsilon_{1},\ldots,\Upsilon_{4}$). We can assume the system is initially in a $\Gamma=+1$ state (this just sets the parity sector for the logical MZMs). Before any lattice surgery measurements, there exist four independent logical operators that stabilize the full system (i.e., the system is in a $+1$ eigenstate of all these operators): $\Gamma$, $\tilde{\Gamma}_{4}$, $i\tilde{\Upsilon}_{1}\tilde{\Upsilon}_{2}\,\bar{X}_{B}$ and $i\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{4}\,\bar{Z}_{B}$. The first measurement operator, $i\Upsilon_{1}\tilde{\Upsilon}_{1}$, anti-commutes with $\Gamma$, $\tilde{\Gamma}_{4}$ and $i\tilde{\Upsilon}_{1}\tilde{\Upsilon}_{2}\,\bar{X}_{B}$, and hence only products of pairs of these three operators continue to stabilize the state of the system. The full list of such (independent) stabilizing logical operators is: $\lambda_{1}i\Upsilon_{1}\tilde{\Upsilon}_{1}$, $\Gamma\tilde{\Gamma}_{4}$, $i\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{4}\,\bar{Z}_{B}$ and $i\tilde{\Upsilon}_{1}\tilde{\Upsilon}_{2}\,\bar{X}_{B}\tilde{\Gamma}_{4}=i\tilde{\Upsilon}_{3}\tilde{\Upsilon}_{4}\,\bar{X}_{B}$. Similarly to Section IV.3.1, we may replace the projector associated with the logical measurement with a logical braid: $\exp(\frac{\pi}{4}\lambda_{1}i\Upsilon_{1}\tilde{\Upsilon}_{1}\tilde{\Gamma}_{4})=\exp(-\frac{\pi}{4}\lambda_{1}i\Upsilon_{1}\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{3}\tilde{\Upsilon}_{4})$. This produces the correct post-measurement state. This operator can be thought of as a braid between $\Upsilon_{1}$ and $i\lambda_{1}\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{3}\tilde{\Upsilon}_{4}$: it maps $\Upsilon_{1}\mapsto i\lambda_{1}\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{3}\tilde{\Upsilon}_{4}$. Thus the operator $\Gamma_{4}$ we wish to measure is mapped to operator $-i\lambda_{1}\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{3}\tilde{\Upsilon}_{4}\Upsilon_{2}\Upsilon_{3}\Upsilon_{4}$.
@@
Note that, similarly to the case of Section IV.3.1, this measurement procedure does not directly affect the bosonic information stored between the twists, as all bosonic logical operators can be deformed away from the locations of the lattice surgery measurements. The introduction of the hole again reduces the distance of some logical qubits, but this distance can be preserved by further separating the twists in the code. Hence all encoded information can remain topologically protected throughout the entire procedure.
IV.4 Magic State Injection
@@
IV.4.1 On-Patch Magic State Injection
Figure 13a shows two fermionic twists (hosting logical MZMs $\Upsilon_{a}$, $\Upsilon_{b}$) at minimum separation. Suppose there exist another two twists at large separation from this pair (hosting logical MZMs $\Upsilon_{c}$, $\Upsilon_{d}$). For this setup, the logical operator $i\gamma_{1}\gamma_{2}$, shown in the figure, equals $i\Upsilon_{a}\Upsilon_{b}\bar{Z}_{B}$, where $\bar{Z}_{B}$ is a logical Pauli operator (green string loop encircling the two logical MZMs, not shown) for the BT qubit stored between the four twists. Hence, $i\gamma_{1}\gamma_{2}$ acts on both FT and BT qubits stored in the twist. To prepare a magic state, we first prepare the state |$\ket{+}_{F}\ket{0}_{B}$ (cf. Sections IV.2 and IV.3.1), where |$\ket{0}_{B}=\bar{Z}_{B}\ket{0}_{B}$ is a computational basis state of the BT qubit and |$\ket{+}_{F}=i\Upsilon_{a}\Upsilon_{c}\ket{+}_{F}$ is a basis state of the FT qubit stored between $\Upsilon_{a},\Upsilon_{b},\Upsilon_{c},\Upsilon_{d}$. We then perform the rotation $\exp{(\frac{\pi}{8}\gamma_{1}\gamma_{2})}$; we assume we can perform at least a noisy version of this gate using established methods [Karzig_geom_magic_2016, Karzig_geom_magic_meas2019, Karzig_Majorana2017]. This approximately yields
@@ IV.4.1 On-Patch Magic State Injection
IV.4.2 Off-Patch Magic State Injection
@@
We start by preparing the patches in the $+1$ eigenstates of $i\Upsilon_{a}\Upsilon_{1}$, $i\Upsilon_{b}\Upsilon_{2}$, $i\Upsilon_{c}\Upsilon_{3}$ and $i\Upsilon_{d}\Upsilon_{4}$ (cf. Section IV.3). We then prepare a patch with four Majorana operators (orange rectangle in Figure 13b and labeled $\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4}$) in a possibly noisy magic state encoded in subspace stabilized by the patch parity $-\gamma_{1}\gamma_{2}\gamma_{3}\gamma_{4}$. We inject this magic state into an FT qubit furnished by $\Upsilon_{a},\ldots,\Upsilon_{d}$ by measuring operators $\mathcal{O}_{j}=i\Upsilon_{j}\gamma_{j}$ for $j=1,2,3,4$. The example of $\mathcal{O}_{4}$ is represented as a plaquette operator in Figure 13b. This measurement maps $\gamma_{1}$ to $\Upsilon_{a}$ (and similarly for the other measurements), by the same logic that we have used repeatedly: due to the anticommutation of $\mathcal{O}_{1}$ and $i\Upsilon_{a}\Upsilon_{1}$, the measurement, with outcome $\lambda_{1}$ effectively implements the logical braid $\exp(\frac{\pi}{4}\lambda_{1}\Upsilon_{a}\gamma_{1})$ and similarly for $j=2,3,4$. In this way, the magic state furnished by $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, $\gamma_{4}$ is transferred to the FT qubit furnished by logical MZMs $\Upsilon_{a},\Upsilon_{b},\Upsilon_{c},\Upsilon_{d}$ (specifically in the subspace stabilized by $-\Upsilon_{a}\Upsilon_{b}\Upsilon_{c}\Upsilon_{d}$), up to a correction operation (if $\lambda_{1}\lambda_{2}\lambda_{3}\lambda_{4}=-1$) which can be tracked. The resource cost of this magic state preparation method is compared with alternatives in Section V.
IV.5 Arbitrary String Operator Measurements
@@
We now discuss measuring various string operators. The following considerations make this particularly interesting. The twist framework established above allows for not only fermionic and bosonic computation to be done side-by-side, but also for hybrid computing schemes involving both FT and BT qubits. To complete the set of universal gates for both types of logical qubit, we require a gate that entangles the two. This would, for example, allow us to use magic states stored in FT qubits (cf. Section IV.4) to apply $T$ gates to BT qubits. Moreover, the measurement of BT qubit logical operators opens up an additional route for implementing Clifford gates on top of braiding twists, which could result in reduced overheads [Lattice_surg_low_overhead].
@@
We can achieve the effect of entangling gates between BT and FT qubits by fault-tolerantly measuring $\Gamma_{F}P_{B}$, where $\Gamma_{F}$ is a logical MZM parity and $P_{B}$ is a bosonic logical Pauli operator. For example, take an FT, a BT, and an ancilla twist qubit, with logical operators $\{\bar{Z}_{F},\bar{X}_{F}\}$, $\{\bar{Z}_{B},\bar{X}_{B}\}$ and $\{\bar{Z}_{A},\bar{X}_{A}\}$, respectively. To implement a CNOT gate controlled on qubit $F$ and targeted on $B$, we can prepare the ancilla qubit in the |$\ket{+}_{A}$ state, then measure $\bar{Z}_{F}\bar{Z}_{A}$, then $\bar{X}_{A}\bar{X}_{B}$, and finally $\bar{Z}_{A}$. This results in the desired gate, up to a logical Pauli operator which can be tracked. We can avoid introducing these ancilla qubits by instead classically tracking Clifford gates and updating circuit measurements accordingly (see also Section V[Bravyi_PBC2016, Tracking_Color_Codes, QC_with_MFCs].
@@
The procedures for measuring string operators use lattice surgery and are very similar to those for measuring fermion parity operators, as described in Section IV.3. We begin with the case of a single fermionic string operator, $S_{1}$. For concreteness, we assume that $S_{1}$ terminates on the boundaries of two (fermionic) $T_{\mathsf{B}}$ twists. As in Section IV.3.1, we move these two twists to a (green- or red-colored) boundary of the lattice, and introduce an ancilla patch of code nearby to this boundary, hosting two logical MZMs $\tilde{\Upsilon}_{a}$ and $\tilde{\Upsilon}_{b}$ (hosted on blue boundaries). Alternatively, a hole can be created in the lattice and the ancilla patch prepared in this hole, as described in previous sections. An example of the resulting setup is shown in Figure 14a. The ancilla patch is initialized in the $+1$ eigenstate of $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}$ (cf. Section IV.3.1), which is a blue string operator. (The red string operator part of $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}$, akin to that in Figure 8a, would now connect the red boundaries and hence can be shrunk: it is a stabilizer element.)
To measure the operator $S_{1}$, we perform lattice surgery between the logical patch and the ancilla patch. For the example of Figure 14, we perform measurements of the green plaquette operators bridging the gap between the two patches. Notice that the product of these green plaquettes is equal to ±$\pm S_{1}(i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b})$. This procedure differs from that of Section IV.3.1 since there we only measure two plaquette operators bridging the gap between the patches. Those plaquette operator measurements fused pairs of logical MZMs on the ancilla and code patches. In the present case we fuse the patches along the two string operators shown in Figure 14 resulting in different logical information being extracted. Call the product of all measurement outcomes of these green plaquette operators $\lambda$. Similarly to the cases examined in Section IV.3, $\lambda$ is the measurement outcome for $S_{1}$ that we desire (up to a known sign defined by the ordering of the operators). To decouple the ancilla patch, we then destructively measure $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}$ (cf. Section IV.3.1). This produces the result $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}$ with certainty (assuming perfect measurements) since we can find a representative of this logical operator that commutes with all lattice surgery measurements. Hence we need not track any correction operation, as we had to do in previously-described lattice surgery procedures.
@@
In Appendix G, we extend these procedures to the measurement of arbitrary products of bosonic and fermionic string operators. Thus, any logical operator in the code can be measured using lattice surgery. We can therefore achieve universal quantum computation with BT and FT qubits side-by-side, by combining these measurements with the procedures of Sections IV.2-IV.4.
@@
V Overheads for Universal Fault-Tolerant Computation

### v4

@@ 4.1 Code Deformation
4.1.1 Moving Bosonic Twists
@@
4.1.2 Moving Fermionic Twists
@@
4.2 State Preparation of BT Qubits and Clifford Gates via Braiding Twists
@@
We will now discuss how to prepare logical Pauli eigenstates and perform fault-tolerant Clifford gates using BT qubits. The latter will be achieved purely by braiding $Z_{2}$ and $Z_{3}$ twists, in analogy to Clifford gate procedures in the bosonic Color Code [Bombin_twists_code_deformation] (though Clifford gates may also be achieved through lattice surgery; see Section 4.5). Our MSC scheme requires no ancilla logical qubits for the implementation of single-qubit Clifford gates, unlike other approaches [MjFerm_Surface_Code, Roadmap_to_MSCs]. We shall consider BT qubits stored sparsely in quartets of either $Z_{2}$ (storing one qubit in four twists) or $Z_{3}$ (two qubits in four twists) bosonic twists. We will see that we need only consider two sub-types of $Z_{2}$ twist, which we take (without loss of generality) to be $T_{\mathsf{B}}$ and $T_{\mathsf{R}}$.
To prepare four $Z_{2}$ twists in an eigenstate of a given logical Pauli operator, we start with a MSC patch hosting these twists, such as the one of Figure 1. For example, to prepare this patch in the $+1$ eigenstate of $\bar{X}$ (a red string operator), we initialize the patch in the $+1$ eigenstate of all red bond operators. This prepares $\bar{X}$ and green and blue plaquette operators in their $+1$ eigenstates since they are all products of red bond operators. To prepare red plaquette operators in their $+1$ eigenstates, we measure them and apply red string operators that take the state to the code space given the observed syndrome (in practice, we would not apply this operator and instead, redefine the stabilizers). Since $\bar{X}$ and the green/blue plaquette operators commute with the red plaquette operators, these measurements do not change their eigenvalues from the preceding step. The inequivalent choices of string in the last step differ by $\bar{X}$; this is again immaterial given we already had a $+1$ eigenstate of $\bar{X}$. This procedure thus prepares the patch in the $+1$ eigenstate of $\bar{X}$ in the code space. We can then move the twists away from the boundaries of the patch using code deformation (cf. Section 4.1). Finally, we can fuse the patch with the rest of the lattice hosting logical information by adding stabilizer operators that bridge the gap between the boundaries of the two lattices.
@@
4.3 State Preparation of FT Qubits and Logical Majorana Parity Measurements via Lattice Surgery
@@
4.3.1 Bilinear Measurement and Eigenstate Preparation
@@
Stabilizing
Logical Operators
@@
4.3.2 Arbitrary Fermion Parity Measurement
@@
To explain the effect of these $i\Upsilon_{j}\tilde{\Upsilon}_{j}$ measurements, we consider the example of Figure 12 more closely. Here we will use $\Gamma$ to denote the total fermion parity operator of the main lattice (containing $\Upsilon_{1},\ldots,\Upsilon_{4}$). We can assume the system is initially in a $\Gamma=+1$ state (this just sets the parity sector for the logical MZMs). Before any lattice surgery measurements, there exist four independent logical operators that stabilize the full system (i.e., the system is in a $+1$ eigenstate of all these operators): $\Gamma$, $\tilde{\Gamma}_{4}$, $i\tilde{\Upsilon}_{1}\tilde{\Upsilon}_{2}\,\bar{X}_{B}$ and $i\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{4}\,\bar{Z}_{B}$. The first measurement operator, $i\Upsilon_{1}\tilde{\Upsilon}_{1}$, anti-commutes with $\Gamma$, $\tilde{\Gamma}_{4}$ and $i\tilde{\Upsilon}_{1}\tilde{\Upsilon}_{2}\,\bar{X}_{B}$, and hence only products of pairs of these three operators continue to stabilize the state of the system. The full list of such (independent) stabilizing logical operators is: $\lambda_{1}i\Upsilon_{1}\tilde{\Upsilon}_{1}$, $\Gamma\tilde{\Gamma}_{4}$, $i\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{4}\,\bar{Z}_{B}$ and $i\tilde{\Upsilon}_{1}\tilde{\Upsilon}_{2}\,\bar{X}_{B}\tilde{\Gamma}_{4}=i\tilde{\Upsilon}_{3}\tilde{\Upsilon}_{4}\,\bar{X}_{B}$. Similarly to Section 4.3.1, we may replace the projector associated with the logical measurement with a logical braid: $\exp(\frac{\pi}{4}\lambda_{1}i\Upsilon_{1}\tilde{\Upsilon}_{1}\tilde{\Gamma}_{4})=\exp(-\frac{\pi}{4}\lambda_{1}i\Upsilon_{1}\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{3}\tilde{\Upsilon}_{4})$. This produces the correct post-measurement state. This operator can be thought of as a braid between $\Upsilon_{1}$ and $i\lambda_{1}\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{3}\tilde{\Upsilon}_{4}$: it maps $\Upsilon_{1}\mapsto i\lambda_{1}\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{3}\tilde{\Upsilon}_{4}$. Thus the operator $\Gamma_{4}$ we wish to measure is mapped to operator $-i\lambda_{1}\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{3}\tilde{\Upsilon}_{4}\Upsilon_{2}\Upsilon_{3}\Upsilon_{4}$.
@@
Note that, similarly to the case of Section 4.3.1, this measurement procedure does not directly affect the bosonic information stored between the twists, as all bosonic logical operators can be deformed away from the locations of the lattice surgery measurements. The introduction of the hole again reduces the distance of some logical qubits, but this distance can be preserved by further separating the twists in the code. Hence all encoded information can remain topologically protected throughout the entire procedure.
4.4 Magic State Injection
@@
4.4.1 On-Patch Magic State Injection
Figure 13a shows two fermionic twists (hosting logical MZMs $\Upsilon_{a}$, $\Upsilon_{b}$) at minimum separation. Suppose there exist another two twists at large separation from this pair (hosting logical MZMs $\Upsilon_{c}$, $\Upsilon_{d}$). For this setup, the logical operator $i\gamma_{1}\gamma_{2}$, shown in the figure, equals $i\Upsilon_{a}\Upsilon_{b}\bar{Z}_{B}$, where $\bar{Z}_{B}$ is a logical Pauli operator (green string loop encircling the two logical MZMs, not shown) for the BT qubit stored between the four twists. Hence, $i\gamma_{1}\gamma_{2}$ acts on both FT and BT qubits stored in the twist. To prepare a magic state, we first prepare the state |$\ket{+}_{F}\ket{0}_{B}$ (cf. Sections 4.2 and 4.3.1), where |$\ket{0}_{B}=\bar{Z}_{B}\ket{0}_{B}$ is a computational basis state of the BT qubit and |$\ket{+}_{F}=i\Upsilon_{a}\Upsilon_{c}\ket{+}_{F}$ is a basis state of the FT qubit stored between $\Upsilon_{a},\Upsilon_{b},\Upsilon_{c},\Upsilon_{d}$. We then perform the rotation $\exp{(\frac{\pi}{8}\gamma_{1}\gamma_{2})}$; we assume we can perform at least a noisy version of this gate using established methods [Karzig_geom_magic_2016, Karzig_geom_magic_meas2019, Karzig_Majorana2017]. This approximately yields
@@ 4.4.1 On-Patch Magic State Injection
4.4.2 Off-Patch Magic State Injection
@@
We start by preparing the patches in the $+1$ eigenstates of $i\Upsilon_{a}\Upsilon_{1}$, $i\Upsilon_{b}\Upsilon_{2}$, $i\Upsilon_{c}\Upsilon_{3}$ and $i\Upsilon_{d}\Upsilon_{4}$ (cf. Section 4.3). We then prepare a patch with four Majorana operators (orange rectangle in Figure 13b and labeled $\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4}$) in a possibly noisy magic state encoded in subspace stabilized by the patch parity $-\gamma_{1}\gamma_{2}\gamma_{3}\gamma_{4}$. We inject this magic state into an FT qubit furnished by $\Upsilon_{a},\ldots,\Upsilon_{d}$ by measuring operators $\mathcal{O}_{j}=i\Upsilon_{j}\gamma_{j}$ for $j=1,2,3,4$. The example of $\mathcal{O}_{4}$ is represented as a plaquette operator in Figure 13b. This measurement maps $\gamma_{1}$ to $\Upsilon_{a}$ (and similarly for the other measurements), by the same logic that we have used repeatedly: due to the anticommutation of $\mathcal{O}_{1}$ and $i\Upsilon_{a}\Upsilon_{1}$, the measurement, with outcome $\lambda_{1}$ effectively implements the logical braid $\exp(\frac{\pi}{4}\lambda_{1}\Upsilon_{a}\gamma_{1})$ and similarly for $j=2,3,4$. In this way, the magic state furnished by $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, $\gamma_{4}$ is transferred to the FT qubit furnished by logical MZMs $\Upsilon_{a},\Upsilon_{b},\Upsilon_{c},\Upsilon_{d}$ (specifically in the subspace stabilized by $-\Upsilon_{a}\Upsilon_{b}\Upsilon_{c}\Upsilon_{d}$), up to a correction operation (if $\lambda_{1}\lambda_{2}\lambda_{3}\lambda_{4}=-1$) which can be tracked. The resource cost of this magic state preparation method is compared with alternatives in Section 5.
4.5 Arbitrary String Operator Measurements
@@
We now discuss measuring various string operators. The following considerations make this particularly interesting. The twist framework established above allows for not only fermionic and bosonic computation to be done side-by-side, but also for hybrid computing schemes involving both FT and BT qubits. To complete the set of universal gates for both types of logical qubit, we require a gate that entangles the two. This would, for example, allow us to use magic states stored in FT qubits (cf. Section 4.4) to apply $T$ gates to BT qubits. Moreover, the measurement of BT qubit logical operators opens up an additional route for implementing Clifford gates on top of braiding twists, which could result in reduced overheads [Lattice_surg_low_overhead].
@@
We can achieve the effect of entangling gates between BT and FT qubits by fault-tolerantly measuring $\Gamma_{F}P_{B}$, where $\Gamma_{F}$ is a logical MZM parity and $P_{B}$ is a bosonic logical Pauli operator. For example, take an FT, a BT, and an ancilla twist qubit, with logical operators $\{\bar{Z}_{F},\bar{X}_{F}\}$, $\{\bar{Z}_{B},\bar{X}_{B}\}$ and $\{\bar{Z}_{A},\bar{X}_{A}\}$, respectively. To implement a CNOT gate controlled on qubit $F$ and targeted on $B$, we can prepare the ancilla qubit in the |$\ket{+}_{A}$ state, then measure $\bar{Z}_{F}\bar{Z}_{A}$, then $\bar{X}_{A}\bar{X}_{B}$, and finally $\bar{Z}_{A}$. This results in the desired gate, up to a logical Pauli operator which can be tracked. We can avoid introducing these ancilla qubits by instead classically tracking Clifford gates and updating circuit measurements accordingly (see also Section 5[Bravyi_PBC2016, Tracking_Color_Codes, QC_with_MFCs].
@@
The procedures for measuring string operators use lattice surgery and are very similar to those for measuring fermion parity operators, as described in Section 4.3. We begin with the case of a single fermionic string operator, $S_{1}$. For concreteness, we assume that $S_{1}$ terminates on the boundaries of two (fermionic) $T_{\mathsf{B}}$ twists. As in Section 4.3.1, we move these two twists to a (green- or red-colored) boundary of the lattice, and introduce an ancilla patch of code nearby to this boundary, hosting two logical MZMs $\tilde{\Upsilon}_{a}$ and $\tilde{\Upsilon}_{b}$ (hosted on blue boundaries). Alternatively, a hole can be created in the lattice and the ancilla patch prepared in this hole, as described in previous sections. An example of the resulting setup is shown in Figure 14a. The ancilla patch is initialized in the $+1$ eigenstate of $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}$ (cf. Section 4.3.1), which is a blue string operator. (The red string operator part of $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}$, akin to that in Figure 8a, would now connect the red boundaries and hence can be shrunk: it is a stabilizer element.)
To measure the operator $S_{1}$, we perform lattice surgery between the logical patch and the ancilla patch. For the example of Figure 14, we perform measurements of the green plaquette operators bridging the gap between the two patches. Notice that the product of these green plaquettes is equal to ±$\pm S_{1}(i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b})$. This procedure differs from that of Section 4.3.1 since there we only measure two plaquette operators bridging the gap between the patches. Those plaquette operator measurements fused pairs of logical MZMs on the ancilla and code patches. In the present case we fuse the patches along the two string operators shown in Figure 14 resulting in different logical information being extracted. Call the product of all measurement outcomes of these green plaquette operators $\lambda$. Similarly to the cases examined in Section 4.3, $\lambda$ is the measurement outcome for $S_{1}$ that we desire (up to a known sign defined by the ordering of the operators). To decouple the ancilla patch, we then destructively measure $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}$ (cf. Section 4.3.1). This produces the result $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}$ with certainty (assuming perfect measurements) since we can find a representative of this logical operator that commutes with all lattice surgery measurements. Hence we need not track any correction operation, as we had to do in previously-described lattice surgery procedures.
@@
In Appendix G, we extend these procedures to the measurement of arbitrary products of bosonic and fermionic string operators. Thus, any logical operator in the code can be measured using lattice surgery. We can therefore achieve universal quantum computation with BT and FT qubits side-by-side, by combining these measurements with the procedures of Sections 4.2-4.4.
@@
5 Overheads for Universal Fault-Tolerant Computation